# Reparo: Loss-Resilient Generative Codec for Video Conferencing

### v2

@@ 4. Evaluation
4.1. Experiment Setup
Baselines. ULPFEC and flexFEC are two solutions included in WebRTC (web, [n. d.]) to recover from audio and video packet loss. Tambur (Rudow et al., 2023) is a recent streaming-codes based FEC solution atop the VP9 video codec (Mukherjee et al., 2015) that has been shown to perform better than classical block-based FEC techniques. Following the original paper, Tambur is implemented in the Ringmaster video conferencing platform (rin, [n. d.]). We set Tambur’s latency deadline to 3 frames, and the bandwidth overhead of all baselines to be approximately 50%. The frame rate is set to and the video resolution to 512512, which is typical for video conferencing.
Datasets. For training the neural codec in Reparo, we use a combination of three datasets: the FFHQ dataset (70,000 images) (Karras et al., 2019), the CelebAHQ dataset (30,000 images) (Karras et al., 2017), and part of the TalkingHeads dataset (25 hours of video) (Wang et al., 2021b). These datasets comprise high-resolution human face images and videos, making them ideal for training Reparo, which is aimed at improving video conferencing quality. We exclusively use the TalkingHeads dataset for training the loss recovery module, as this module operates on video clips instead of images.
For evaluating Reparo and the baselines in the context of video conferencing applications, we combine the dataset from Gemino (Sivaraman et al., 2022) and part of the TalkingHeads dataset (Wang et al., 2021b), forming a large and diverse video-conferencing validation dataset consisting of 5 hours of data from 412 video clips and 84 different subjects. We make sure that there is no overlap between the training and the validation set. Notably, our validation set is much larger and more diverse than prior works on video conferencing, such as Gemino (Sivaraman et al., 2022) (25 videos from 5 people, 75 minutes in total), Tambur (Rudow et al., 2023) (20 videos, 200 minutes in total), and GRACE (Cheng et al., 2023) (60 videos, 745 seconds in total), demonstrating the generalization ability and effectiveness of Reparo.
Implementation. We studied two FEC algorithms, ULPFEC and flexFEC, implemented in Google’s implementation of WebRTC (web, [n. d.]) with VP8 codec. The FEC rate is set to 50% with packet retransmissions disabled, and the frame wait-time is set to . This wait-time parameter dictates the duration after which WebRTC stops to decode a frame that has been partially received. Our experimental setup includes a headless real-time video application, built on the WebRTC platform, which processes video input at a rate of at the sender’s end and records the output to a file at the receiver’s end. The network conditions between the sender and receiver were simulated using the Mahimahi network emulator (Netravali et al., 2015). Within this environment, we constructed a loss shell akin to that used in Tambur, facilitating controlled network loss scenarios for our experiments. Tambur is evaluated using the Ringmaster platform with VP9 codec.
@@ 4.1. Experiment Setup
Metrics. We evaluate multiple metrics, including peak signal-to-noise ratio (PSNR), structural similarity (SSIM), Learned Perceptual Image Patch Similarity (LPIPS), and percentage of non-rendered frames. PSNR, SSIM, and LPIPS are computed by comparing the displayed videos at the receiver to the original videos at the transmitter, using the PyTorch Image Quality (PIQ) library (Kastryulin et al., 2022). Non-rendered frames are defined differently for baselines and Reparo. For baselines, we compute the percentage of frames that are not played by the receiver due to packet loss on that frame itself or dependency on previously undecodable frames. For Reparo, we define “non-rendered frames” as those frames with PSNR less than , as our scheme always tries to generate and render a frame. We have observed that for our dataset, VP8/9’s PSNR rarely drops below unless there’s an undecodable frame, so treating Reparo’s frames with PSNR below as “non-rendered” would favor baselines in comparisons. Non-rendered frames correlate well with standard quality-of-experience (QoE) metrics, as a large number of non-rendered frames can lead to video freezes and degrade QoE. All metrics are aggregated over all frames of the videos in our validation dataset. We compute bitrate by averaging the packet sizes (without TCP/IP headers) recorded in the network traces over the course of the entire video.
Network Scenarios. We consider two primary scenarios that can result in packet loss during transmission: (1) an unreliable network (§4.2), and (2) a rate-limited link (§4.3). In the first scenario, the network is unreliable and randomly drops packets due to poor conditions. To simulate this scenario, we use a GE loss channel that transitions between a “good” state with a low packet loss rate and a “bad” state with a high packet loss rate, similar to Tambur’s setup (Rudow et al., 2023). The probability of transitioning from the good state to the bad state and vice versa is 0.068 and 0.852, respectively. The probability of loss in the good state is 0.04. These parameters are set to mimic Tambur’s evaluation, and are computed to approximate the actual statistics over a large corpus of traces from Microsoft Teams (Rudow et al., 2023). We vary the probability of loss in the bad state to evaluate baselines and Reparo’s performance under different loss levels. Specifically, we set it to 0.25 to simulate a low loss level, 0.5 to simulate a medium loss level, and 0.75 to simulate a high loss level (the default value in Tambur’s evaluation is 0.5). The statistics of the simulated bursty loss networks are included in Tab. 1. In the second scenario, we consider a fixed-rate link that drops packets once saturated. To simulate this, we use a FIFO queue with a fixed queue length of 6KB and a drain rate of 320 Kbps.
@@
4.2. Performance on Lossy Networks
@@
We further show the distribution of frame PSNR values across the frames in our evaluation corpus with Reparo and baselines at under different packet loss rates in the “bad” state of the GE channel. As shown in Fig. 6, Reparo’s distribution and averages of PSNR values are more or less unaffected by the loss level. With Reparo, almost 99% of frames have PSNR values larger than . The variance of PSNR values across displayed frames is also much lower with Reparo than the baselines, showing the stability of the quality of the displayed video across loss levels. Specifically, Reparo’s frame PSNR values are mostly between and (¿90%)). In contrast, as the loss level becomes higher, the baselines are more likely to experience video freezes, resulting in more frames with low PSNR. These results demonstrate that Reparo is more robust and efficient at recovering from packet losses than current FEC schemes for video conferencing.
@@
4.3. Performance on Rate-Limited Networks
@@
Such a stable bitrate improves Reparo’s performance over fixed-capacity bottleneck links. To simulate such a link, we use a FIFO queue with a constant (drain) rate of Kbps. The size of the queue is set to , as such a queue will introduce a delay, which is the upper bound of industry recommendations for interactive video conferencing (Union, 2001). Packets are queued first and drained at the desired link rate. When the FIFO queue becomes full, subsequent packets will be dropped. In Fig. 11, we set to and show the average PSNR of Reparo and VP9+Tambur with different target bitrates for each codec. Note that the target bitrate for VP9+Tambur typically does not match the actual bitrate: it is the input parameter for the VP9 codec to encode a video. As a result, the actual bitrate of VP9+Tambur can be much larger than the target bitrate of VP9 depending on the encoding speed and quality parameters. Also, Tambur’s parity packets typically introduce 50% to 60% bandwidth overhead, further inflating the actual bitrate of VP9+Tambur. For example, the target bitrate corresponds to an actual average bitrate of . As a result, we only vary the target bitrate supplied to VP9+Tambur up to because beyond that its actual bitrate with FEC overheads overshoots the link rate and causes a lot of packet drops. On the other hand, Reparo’s actual bitrate can exactly match the target bitrate.
@@
In practice, congestion control protocols like GCC (Carlucci et al., 2016) are used to adapt the encoder’s target bitrate based on network observations such as latency and loss. However, this experiment shows that choosing the appropriate target bitrate for Tambur is much more challenging than for Reparo. For VP9+Tambur, the adaptation protocol must be conservative and operate in a lower bitrate range to limit packet drops. In contrast, Reparo can continue to benefit from larger target bitrates as long as they are smaller than the link capacity, and the best performance is achieved by setting the target bitrate near the link capacity.
4.4. Other Results
@@
Latency. In Tab. 2, we present the latency of different modules in Reparo. The neural codec and loss recovery modules of Reparo have higher encoding and decoding latencies compared to traditional FEC schemes, since they require heavy computation. For our implementation, the total inference delay incurred by Reparo is . With typical network queuing delays of , the end-to-end delay of Reparo is less than , which meets the industry recommendation of for maximum tolerable latency for interactive video applications (Union, 2001).
@@
5. Limitations
Although Reparo offers several key advantages over past work, it also has some limitations. First, the current implementation of Reparo is in PyTorch, and uses transformers which are computationally more intensive than traditional video codecs and FEC-based methods (Dosovitskiy et al., 2021; Arnab et al., 2021). It requires GPUs equivalent to an Apple M2 Max GPU to operate in real time. This limits the range of devices on which Reparo can be deployed, and the current implementation is not suitable for low-end devices such as smartphones or tablets. However, machine learning models can be sped up for edge devices using more efficient model architectures (Howard et al., 2017; Tan and Le, 2019; Elsken et al., 2019; Rao et al., 2021), hardware design (Gale et al., 2020; Han et al., 2017; Zhang et al., 2020), and techniques such as knowledge distillation (Hinton et al., 2015). We leave an investigation of such optimizations to future work. We also note that over time more powerful GPUs are integrated in edge devices naturally paving the way for running complex neural networks on them. Second, Reparo learns a dictionary of tokens specific to the domain of interest, which is video conferencing for the purpose of this paper. While Reparo can potentially be extended beyond video conferencing to other domains, this will require learning a dictionary for each new domain. As such it is more specialized and less general than traditional FEC codecs.
@@ Appendix A Implementation Details
Because of GPU memory limitation, we adopt a two-stage training recipe for Reparo similar to many prior approaches (Li et al., 2022; Esser et al., 2021). We first train our VQGAN codec which encodes each video frame into discrete tokens without any losses. We then fix the VQGAN codec and train the loss recovery module on the discrete tokens with self-dropping and packet loss. In this section, we describe the neural network structure and training schemes of Reparo’s neural codec and loss recovery module, as well as the design of the bitrate controller in detail.
A.1. Neural Codec
@@
Model Structure. We use a CNN-based VQGAN (Esser et al., 2021) encoder and quantizer to tokenize the input frame to quantized features, where 128 is the number of channels of the quantized features. It then uses a codebook to quantize the features by finding the nearest neighbor of each feature in the codebook. The codebook is a matrix by default, with 1024 entries, each of which uses a 128-dimensional feature. The decoder operates on the quantized features and reconstructs the video frame. The encoder consists of 5 blocks and each block consists of 2 residual blocks which follow standard ResNet’s residual block design (He et al., 2016). After each block in the encoder, the feature vector is down-sampled by 2 using average pooling. The quantizer then maps each pixel of the encoder’s output feature map to the nearest token (based on distance) in the codebook with entries, each entry with channels. The decoder consists of another 5 blocks where each encoder block has 2 residual blocks. After each block in the decoder, the feature map is up-sampled by 2 using bicubic interpolation. The tokenizer consists of 23.8M parameters and the detokenizer consists of 30.5M parameters.
@@
Training schemes. We follow the original VQGAN training recipe (Esser et al., 2021) to train the VQGAN. We use a vector-quantize loss between the image features and quantized tokens that nudges the image features towards the tokens that they map to, a reconstruction loss () between the input and final reconstructed frame, a perceptual loss (Zhang et al., 2018) between the input and reconstructed frame, and a discriminative loss (Isola et al., 2017) between the input and reconstructed frame. Detailed descriptions of the losses can be found in the VQGAN paper (Esser et al., 2021).
@@
We use the officially released VQGAN encoder and decoder pre-trained on OpenImages (Kuznetsova et al., 2020) to initialize our codec whenever possible. OpenImages is a large-scale image dataset consisting of 9M natural images. We observe that such initialization largely speeds up our training (takes 10 epochs to converge), but we also note that training from scratch on our pre-training face datasets can achieve similar performance with a much longer training time (200 epochs). We train our neural codec using a constant learning rate and train it until there is no substantial change in the training loss. Please refer to Tab. 3 for the training recipe of our VQGAN codec.
A.2. Loss Recovery Module
@@
Model Structure. The major component of our loss recovery module is a spatio-temporal ViT network. In our default setting, the input tokens are of shape , where , , , . We use two separate learnable position embeddings, one for time and one for space, which we add together to provide each input token its positional information. We then adopt a standard spatio-temporal ViT architecture (Dosovitskiy et al., 2021), which consists of a stack of spatio-temporal Transformer blocks (Vaswani et al., 2017). Each spatio-temporal block consists of a spatial block and a temporal block. Each of the two blocks independently consists of a multi-head self-attention block and a multi-layer perceptron (MLP) block. In total, we use 20 spatio-temporal Transformer blocks. The number of heads in each multi-head self-attention layer is 12, and the MLP ratio is 4. The embedding dimension throughout the Transformer is 768. Our spatio-temporal ViT consists of 172M parameters. We note that more Transformer blocks, more heads in the self-attention layer, and a larger embedding dimension can further improve the performance of Reparo , but they also introduce more computation overheads.
@@
| Optimizer | Adam (Loshchilov and Hutter, 2017) |
@@
| Learning Rate Schedule | Cosine Decay (Loshchilov and Hutter, 2016) |

### v3

@@ 4 Evaluation
4.1 Experiment Setup
Baselines. ULPFEC and flexFEC are two solutions included in WebRTC [3] to recover from audio and video packet loss. Tambur [44] is a recent streaming-codes based FEC solution atop the VP9 video codec [38] that has been shown to perform better than classical block-based FEC techniques. Following the original paper, Tambur is implemented in the Ringmaster video conferencing platform [2]. We set Tambur’s latency deadline to 3 frames, and the bandwidth overhead of all baselines to be approximately 50%. The frame rate is set to and the video resolution to 512512, which is typical for video conferencing.
Datasets. For training the neural codec in Reparo, we use a combination of three datasets: the FFHQ dataset (70,000 images) [26], the CelebAHQ dataset (30,000 images) [25], and part of the TalkingHeads dataset (25 hours of video) [54]. These datasets comprise high-resolution human face images and videos, making them ideal for training Reparo, which is aimed at improving video conferencing quality. We exclusively use the TalkingHeads dataset for training the loss recovery module, as this module operates on video clips instead of images.
For evaluating Reparo and the baselines in the context of video conferencing applications, we combine the dataset from Gemino [46] and part of the TalkingHeads dataset [54], forming a large and diverse video conferencing validation dataset consisting of 5 hours of data from 412 video clips and 84 different subjects. We make sure that there is no overlap between the training and the validation set. Notably, our validation set is much larger and more diverse than prior works on video conferencing, such as Gemino [46] (25 videos from 5 people, 75 minutes in total), Tambur [44] (20 videos, 200 minutes in total), and GRACE [11] (60 videos, 745 seconds in total), demonstrating the generalization ability and effectiveness of Reparo.
Implementation. We studied two FEC algorithms, ULPFEC and flexFEC, implemented in Google’s implementation of WebRTC [3] with VP8 codec. The FEC rate is set to 50% with packet retransmissions disabled, and the frame wait-time is set to . This wait-time parameter dictates the duration after which WebRTC stops to decode a frame that has been partially received. Our experimental setup includes a headless real-time video application, built on the WebRTC platform, which processes video input at a rate of at the sender’s end and records the output to a file at the receiver’s end. The network conditions between the sender and receiver were simulated using the Mahimahi network emulator [39]. Within this environment, we constructed a loss shell akin to that used in Tambur, facilitating controlled network loss scenarios for our experiments. Tambur is evaluated using the Ringmaster platform with VP9 codec.
@@ 4.1 Experiment Setup
Metrics. We evaluate multiple metrics, including peak signal-to-noise ratio (PSNR), structural similarity (SSIM) [55], Learned Perceptual Image Patch Similarity (LPIPS) [58], and percentage of non-rendered frames. PSNR, SSIM, and LPIPS are computed by comparing the displayed videos at the receiver to the original videos at the transmitter, using the PyTorch Image Quality (PIQ) library [27]. Non-rendered frames are defined differently for baselines and Reparo. For baselines, we compute the percentage of frames that are not played by the receiver due to packet loss on that frame itself or dependency on previously undecodable frames. For Reparo, we define “non-rendered frames” as those frames with PSNR less than , as our scheme always tries to generate and render a frame. We have observed that for our dataset, VP8/9’s PSNR rarely drops below unless there’s an undecodable frame, so treating Reparo’s frames with PSNR below as “non-rendered” would favor baselines in comparisons. Non-rendered frames correlate well with standard quality-of-experience (QoE) metrics, as a large number of non-rendered frames can lead to video freezes and degrade QoE. All metrics are aggregated over all frames of the videos in our validation dataset. We compute bitrate by averaging the packet sizes (without TCP/IP headers) recorded in the network traces over the course of the entire video.
Network Scenarios. We consider two primary scenarios that can result in packet loss during transmission: (1) an unreliable network (§4.2), and (2) a rate-limited link (§4.3). In the first scenario, the network is unreliable and randomly drops packets due to poor conditions. To simulate this scenario, we use a GE loss channel. This channel transitions between a “good” state with a low packet loss rate and a “bad” state with a high packet loss rate, similar to Tambur’s setup [44]. The probability of transitioning from the good state to the bad state and vice versa is 0.068 and 0.852, respectively. The probability of loss in the good state is 0.04. These parameters are set to mimic Tambur’s evaluation, and are computed to approximate the actual statistics over a large corpus of traces from Microsoft Teams [44]. We vary the probability of loss in the bad state to evaluate baselines and Reparo’s performance under different loss levels. Specifically, we set it to 0.25 to simulate a low loss level, 0.5 to simulate a medium loss level, and 0.75 to simulate a high loss level (the default value in Tambur’s evaluation is 0.5). The statistics of the simulated bursty loss networks are included in Tab. 1. In the second scenario, we consider a fixed-rate link that drops packets once saturated. To simulate this, we use a FIFO queue with a fixed queue length of 6KB and a drain rate of 320 Kbps.
@@
4.2 Performance on Lossy Networks
@@
We further show the distribution of frame PSNR values across the frames in our evaluation corpus with Reparo and baselines at under different packet loss rates in the “bad” state of the GE channel. As shown in Fig. 6, Reparo’s distribution and averages of PSNR values are more or less unaffected by the loss level. With Reparo, almost 99% of frames have PSNR values larger than . The variance of PSNR values across displayed frames is also much lower with Reparo than the baselines, showing the stability of the quality of the displayed video across loss levels. Specifically, Reparo’s frame PSNR values are mostly between and (>90%)). In contrast, as the loss level becomes higher, the baselines are more likely to experience video freezes, resulting in more frames with low PSNR. These results demonstrate that Reparo is more robust and efficient at recovering from packet losses than current FEC schemes for video conferencing.
@@
4.3 Performance on Rate-Limited Networks
@@
Such a stable bitrate improves Reparo’s performance over fixed-capacity bottleneck links. To simulate such a link, we use a FIFO queue with a constant (drain) rate of Kbps. The size of the queue is set to , as such a queue will introduce a delay, which is the upper bound of industry recommendations for interactive video conferencing [50]. Packets are queued first and drained at the desired link rate. When the FIFO queue becomes full, subsequent packets will be dropped. In Fig. 11, we set to and show the average PSNR of Reparo and VP9+Tambur with different target bitrates for each codec. Note that the target bitrate for VP9+Tambur typically does not match the actual bitrate: it is the input parameter for the VP9 codec to encode a video. As a result, the actual bitrate of VP9+Tambur can be much larger than the target bitrate of VP9 depending on the encoding speed and quality parameters. Also, Tambur’s parity packets typically introduce 50% to 60% bandwidth overhead, further inflating the actual bitrate of VP9+Tambur. For example, the target bitrate corresponds to an actual average bitrate of . As a result, we only vary the target bitrate supplied to VP9+Tambur up to because beyond that its actual bitrate with FEC overheads overshoots the link rate and causes a lot of packet drops. On the other hand, Reparo’s actual bitrate can exactly match the target bitrate.
@@
In practice, congestion control protocols like GCC [8] are used to adapt the encoder’s target bitrate based on network observations such as latency and loss. However, this experiment shows that choosing the appropriate target bitrate for Tambur is much more challenging than for Reparo. For VP9+Tambur, the adaptation protocol must be conservative and operate in a lower bitrate range to limit packet drops. In contrast, Reparo can continue to benefit from larger target bitrates as long as they are smaller than the link capacity, and the best performance is achieved by setting the target bitrate near the link capacity.
4.4 Other Results
@@
Latency. In Tab. 2, we present the latency of different modules in Reparo. The neural codec and loss recovery modules of Reparo have higher encoding and decoding latencies compared to traditional FEC schemes, since they require heavy computation. For our implementation, the total inference delay incurred by Reparo is . With typical network queuing delays of , the end-to-end delay of Reparo is less than , which meets the industry recommendation of for maximum tolerable latency for interactive video applications [50].
@@
5 Limitations
Although Reparo offers several key advantages over past work, it also has some limitations. First, the current implementation of Reparo is in PyTorch, and uses transformers which are computationally more intensive than traditional video codecs and FEC-based methods [14, 4]. It requires GPUs equivalent to an Apple M2 Max GPU to operate in real time. This limits the range of devices on which Reparo can be deployed, and the current implementation is not suitable for low-end devices such as smartphones or tablets. However, machine learning models can be sped up for edge devices using more efficient model architectures [22, 49, 15, 40], hardware design [17, 18, 59], and techniques such as knowledge distillation [21]. We leave an investigation of such optimizations to future work. We also note that over time more powerful GPUs are integrated in edge devices naturally paving the way for running complex neural networks on them. Second, Reparo learns a dictionary of tokens specific to the domain of interest, which is video conferencing for the purpose of this paper. While Reparo can potentially be extended beyond video conferencing to other domains, this will require learning a dictionary for each new domain. As such it is more specialized and less general than traditional FEC codecs.
@@ Appendix A Implementation Details
Because of GPU memory limitation, we adopt a two-stage training recipe for Reparo similar to many prior approaches [31, 16]. We first train our VQGAN codec which encodes each video frame into discrete tokens without any losses. We then fix the VQGAN codec and train the loss recovery module on the discrete tokens with self-dropping and packet loss. In this section, we describe the neural network structure and training schemes of Reparo’s neural codec and loss recovery module, as well as the design of the bitrate controller in detail.
A.1 Neural Codec
@@
Model Structure. We use a CNN-based VQGAN [16] encoder and quantizer to tokenize the input frame to quantized features, where 128 is the number of channels of the quantized features. It then uses a codebook to quantize the features by finding the nearest neighbor of each feature in the codebook. The codebook is a matrix by default, with 1024 entries, each of which uses a 128-dimensional feature. The decoder operates on the quantized features and reconstructs the video frame. The encoder consists of 5 blocks and each block consists of 2 residual blocks which follow standard ResNet’s residual block design [20]. After each block in the encoder, the feature vector is down-sampled by 2 using average pooling. The quantizer then maps each pixel of the encoder’s output feature map to the nearest token (based on distance) in the codebook with entries, each entry with channels. The decoder consists of another 5 blocks where each encoder block has 2 residual blocks. After each block in the decoder, the feature map is up-sampled by 2 using bicubic interpolation. The tokenizer consists of 23.8M parameters and the detokenizer consists of 30.5M parameters.
@@
Training schemes. We follow the original VQGAN training recipe [16] to train the VQGAN. We use a vector-quantize loss between the image features and quantized tokens that nudges the image features towards the tokens that they map to, a reconstruction loss () between the input and final reconstructed frame, a perceptual loss [58] between the input and reconstructed frame, and a discriminative loss [23] between the input and reconstructed frame. Detailed descriptions of the losses can be found in the VQGAN paper [16].
@@
We use the officially released VQGAN encoder and decoder pre-trained on OpenImages [29] to initialize our codec whenever possible. OpenImages is a large-scale image dataset consisting of 9M natural images. We observe that such initialization largely speeds up our training (takes 10 epochs to converge), but we also note that training from scratch on our pre-training face datasets can achieve similar performance with a much longer training time (200 epochs). We train our neural codec using a constant learning rate and train it until there is no substantial change in the training loss. Please refer to Tab. 3 for the training recipe of our VQGAN codec.
A.2 Loss Recovery Module
@@
Model Structure. The major component of our loss recovery module is a spatio-temporal ViT network. In our default setting, the input tokens are of shape , where , , , . We use two separate learnable position embeddings, one for time and one for space, which we add together to provide each input token its positional information. We then adopt a standard spatio-temporal ViT architecture [14], which consists of a stack of spatio-temporal Transformer blocks [52]. Each spatio-temporal block consists of a spatial block and a temporal block. Each of the two blocks independently consists of a multi-head self-attention block and a multi-layer perceptron (MLP) block. In total, we use 20 spatio-temporal Transformer blocks. The number of heads in each multi-head self-attention layer is 12, and the MLP ratio is 4. The embedding dimension throughout the Transformer is 768. Our spatio-temporal ViT consists of 172M parameters. We note that more Transformer blocks, more heads in the self-attention layer, and a larger embedding dimension can further improve the performance of Reparo , but they also introduce more computation overheads.
@@
| Optimizer | Adam [35] |
@@
| Learning Rate Schedule | Cosine Decay [34] |